As the two events are independent multiply the respective probabilities together:
pr(5) = 1/6
pr(tails) = 1/2*
→ pr(5 and tails) = 1/6 × 1/2 = 1/12
*Actually it is slightly less than 1/2 as there is a very small probability that the coin will end up on its edge; for a small coin it is so small that it can be ignored, but with a scaled up coin it becomes a real possibility as (accidentally) demonstrated in the 1997 Royal Institution Christmas Lecture "The Magical Maze"
Assuming you want the probability FOR A SINGLE TRY, and you want the numbers in that exact order, the probability for each part (for instance, first = red; or second = green) is 1/4; therefore, the probability for the combination is (1/4) to the power 4.
The chance that a given event will occur usually expressed between the number 0 will not and 1 the event will occur
What is the probability of pulling a 5 or a heart from a deck of cards?
Well, if you include the jokers, you have 52 cards total. There are four suits, so that's four "5" cards in the deck, which gives you a 4 in 52 (or 2 in 21, simplified) probability. Each suit has 13 cards, so you have a 13 in 52 chance of pulling a card of any particular suit (13 is a prime number, and can't be simplified).
It is 1/13*1/51 = 1/663 = 0.001508 approx.
25.00
What is the probability of getting a heart out of a pack of cards?
If the pick is completely random, the deck is a standard deck and there are no jokers or any other cards other than the standard 52, the probability is 1/4
What does data distribution mean?
It is the set of values that a variable can take together with the probability or frequency distribution for those values.
How many simple random samples of size 5 can be selected from a population of size 35?
There are 324,632 possible samples.
P(R) = 0.7
The correct choice of answer is B. 71%.
What is Variance in normal distribution?
This browser is not much use when it comes to mathematics but I'll try.
Suppose X is a random variable with a Normal distribution and let f(x) be the probability density function of x.
Then the mean is mu = E(X) = Integral of x*f(x) dx over the domain of X [which is negative infinity to positive infinity].
The variance is E{[X - E(X)]2} = Integral of (x - mu)2*f(x) dx over the domain of X.
What is the probability of drawing a 6-inch pair of pliers from the first basket without looking?
To answer that we need to know how many baskets there are and how many of the different things there are in each basket.
Else - If there were only pliers in the first basket OR there was only one set of pliers and nothing else then the probility is at 100% that you'll get them on first seclection
What data cant be measured but divided into groups?
I suspect that the answer is meant to be qualitative data but that is not a proper answer. Information about the qualitative aspect of data (eg what colour is you hair) is still a measurement. It may not be numerical measurement, but the question states "can't be measured", not "can't be measured numerically".
44%
If you flip a coin what is the chance of getting tails?
A coin has two (major) sides and neither side has any more advantage than the other.
100 (percent) divided by 2 (sides) = 50
There is a 50% chance (or 1/2) of the coin landing on tails.
7*6*5*4*3*2*1= 5040 different ways
The probability is 0.5
What is the probability of observing a sample mean of 18 or more from a sample size of 35?
Without knowing the data which is being sampled, it is impossible to answer this other than by saying that the probability is between 0 and 1 inclusive.
Consider a company. If you sample the annual pay of the employees, any mean will be greater than 18 as everyone will be taking home more than £18 per year, so the probability is 1.
Consider a school. If you sample the lengths of feet of the pupils, any mean will be less than 18 as all the feet are less than 18 inches long, so the probability is 0.
Simply count the prime numbers less than or equal to 50. Then divide that by 50.